A metallic table tennis ball of mass 0.127 g has a charge of 7.3E-6 C.  What potential difference is required to hold the table tennis ball in suspension across a large parallel plate apparatus with plate separation of 20.0 cm.


1
Expert's answer
2021-05-31T08:39:17-0400

In order to hold the table tennis ball in suspension across a large parallel plate apparatus the force of gravity acting on the ball must be balanced by the electric force:


"F_e=F_g,""qE=mg,""\dfrac{q\Delta V}{d}=mg,""\Delta V=\dfrac{mgd}{q},""\Delta V=\dfrac{1.27\cdot10^{-4}\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot0.2\ m}{7.3\cdot10^{-6}\ C}=34\ V."